# Detecting Latin in Historical Books with Large Language Models: A Multimodal Benchmark

### v2

@@ Abstract
This paper presents a novel task of extracting Latin fragments from mixed-language historical documents with varied layouts. We benchmark and evaluate the performance of large foundation models against a multimodal dataset of 724 annotated pages. The results demonstrate that reliable Latin detection with contemporary models is achievable. Our study provides the first comprehensive analysis of these models’ capabilities and limits for this task.  [^1] Under review. Both the dataset and code will be published.
@@ 2 Problem Definition
Formally, given a document page $D$, let $I_{D}$ denote its image and $T_{D}$ denote its OCR’d text. A system must perform the following two subtasks:
@@
We structure our problem into two tasks for a more comprehensive evaluation. The core challenge, Task 2, is the fine-grained extraction of Latin text segments, evaluated using per-page token precision and recall. Since these metrics inadequately handle pages with no Latin, we introduce Task 1, a page-level binary classification, to assess performance on these negative instances specifically. We require extracted segments as strings rather than image regions, as this output format better aligns with the capabilities of most MLLMs and enables a simpler, more direct comparison across different input modalities. Detailed definitions of our metrics are provided in Section 5.
@@ 3 Related work
Given its historical importance, Latin has attracted considerable attention within the NLP community (e.g., Sprugnoli et al., 2024; Schulz and Keller, 2016; Sprugnoli et al., 2022; Gorovaia et al., 2024; Perrone et al., 2021; Burns et al., 2021), though much of this research has centered on small, clean corpora of ancient literary texts. While some recent studies have ventured into Early Modern mixed-language documents (Stüssi and Ströbel, 2024; Volk et al., 2024), these also predominantly rely on manually curated and annotated data. In contrast, our work focuses on the foundational task of Latin discovery: detecting Latin within extensive, unedited, and noisy digitized collections like ECCO Tolonen et al. (2022). This computational approach aims to detect Latin in a vast corpora, while the identified fragments can subsequently be analyzed using a range of established NLP tools developed for classical languages (Johnson et al., 2021; Burns, 2023; Straka and Straková, 2020; Kupari et al., 2024).
@@ 4.1 Sampling and Annotations
These pages were then manually annotated. Annotators were tasked with drawing bounding boxes around all Latin fragments on the page images (see example in Figure 1). These visual annotations later are reliably mapped to text offsets (locations within a string) using ECCO’s OCR positional data for ground truth text extraction. Our annotation guidelines stated marking all instances of Latin text semantically used as Latin. This included single Latin words if presented with explanations in a dictionary, as well as Latin found in headlines, editorial annotations, or footnotes.
The annotation environment was Label Studio Tkachenko et al. (2020-2025). The primary annotation was performed by three scholars familiar with Latin. Following this, an expert in historical texts meticulously reviewed and validated all annotations to ensure accuracy and consistency. Afterwards, the annotators veto the pages with incompletely OCR-transcribed regions.
@@ 12. Side-note 8
Table 1: Page frequencies with categorized segment.
@@ 5.1 OCR Post-Correction and Normalization
Even after the post-correction, residual noise and other variation still remain in the data. Thus, for token-based evaluation, we apply a more traditional rule-based pre-processing pipeline to both predicted and reference strings. This deterministic pipeline, informed by our extensive experience with OCR data and domain-specific knowledge, targets common superficial textual variations. The pipeline includes Unicode normalization, ligature replacement, lowercasing, digit removal, de-hyphenation, and punctuation stripping. Subsequent to these cleaning operations, the strings are tokenized into word-level units. More details on the processing steps are presented in Appendix B.1.
@@ 5.2 Metrics
The goal of Task 1 is to detect whether a page has Latin on it. We measure this by reporting precision, recall, and F1-score in percentage. To evaluate the Latin segment extraction performance in Task 2, we calculate precision, recall, and F1 score in percentage based on token-level matches between model predictions and the ground truth of the page. A fuzzy matching mechanism is applied to pair predicted and reference tokens one by one. A match is considered valid if the token-level edit-distance is not larger than a tunable threshold proportion $\theta$ compared to the ground-truth token length. This approach provides a more flexible and robust evaluation than exact token matching by tolerating minor textual differences at the token level, such as lexical variations and OCR-induced distortions. The pseudocode of the fuzzy matching algorithm is shown in the Appendix B.2. Overall metrics are averaged across the full evaluation set.
@@ 7.1 Experiment Setup
Model Details Page-level (Task 1) Token-level (Task 2) Model Variant Size Mode F1 Precision Recall F1 Precision Recall Lingua - - T 92.58 86.32 99.83 77.14 77.31 80.07 GPT-4.1 - - I+T 96.03 92.51 99.83 84.59 88.49 83.54 Deepseek-R1 - 671B T 97.53 95.34 99.83 87.00 89.69 86.17 Distill-Llama 70B T 96.18 92.94 99.66 84.26 87.81 83.62 Distill-Qwen 32B T 95.35 92.55 98.32 82.84 86.14 82.66 Qwen3 MoE-A22B 235B T 96.90 94.13 99.83 86.54 89.16 86.00 MoE-A3B 30B T 95.90 93.45 98.48 84.29 87.29 83.73 - 32B T 95.39 93.25 97.64 84.07 86.46 83.98 - 14B T 96.61 94.96 98.32 84.78 88.10 83.71 - 4B T 88.45 87.36 89.56 75.38 76.30 77.20 Qwen2.5 VL 72B I+T 98.58 98.16 98.99 83.87 86.98 82.74 VL 72B I 98.82 98.66 98.99 80.00 82.36 79.72 - 72B T 97.85 100.00 95.79 80.95 85.31 79.33 VL 32B I+T 96.18 92.94 99.66 84.32 86.90 83.99 VL 32B I 96.80 94.40 99.33 79.82 82.64 79.18 - 32B T 97.26 99.13 95.45 81.57 84.46 80.99 VL 7B I+T 88.36 82.89 94.61 72.41 78.62 72.23 VL 7B I 95.75 96.74 94.78 71.42 76.99 70.39 - 7B T 91.86 92.49 91.25 60.52 65.78 64.57 InternVL3 - 38B I+T 95.62 92.32 99.16 84.24 84.43 86.67 - 38B I 89.51 87.22 91.92 56.53 61.11 55.61 - 14B I+T 94.28 90.42 98.48 81.36 82.63 83.87 - 14B I 90.16 87.86 92.59 53.51 56.46 55.08 - 8B I+T 90.54 86.39 95.12 69.04 65.87 82.10 - 8B I 90.82 88.50 93.27 60.22 60.30 65.63 Gemma3 - 27B I+T 90.77 83.57 99.33 82.50 84.00 84.79 - 27B I 88.42 82.60 95.12 60.03 62.78 61.68 - 27B T 94.22 90.03 98.82 83.79 86.09 84.20
Table 2: Experimental results on selected LLMs, compared with Lingua baseline. “I” indicates image input, “T” indicates text input, and “I + T” refers to the combination of both modalities.
7.2 Model Results and Comparisons
@@ 7.3 Performance by Category
7.4 Prompt Modification Experiments
Figure 3: Prompt modification experiment results.
To assess the impact of prompt engineering, we evaluated the top mid-size multimodal Qwen2.5-VL-32B model with various strategies, shown as Figure 3. Most strategies result in only marginal changes, typically a trade-off between precision and recall, or different tasks, rather than a definitive improvement. We highlight the most salient strategies here, with full tables of all the prompts and results available in Appendix E.2. For instance, instructing the model to include “single-word” segments lowered page-level precision by encouraging over-extraction on short segments, while negative constraints like “no borrow” slightly increased token-level precision through a more conservative behavior by excluding the borrowed Latin words in other languages. Most revealingly, the “partial categories” strategy, directing the model to focus on its known weak areas, increased token recall but at the cost of a drop in precision. This indicates the prompt did not grant the model a deeper functional understanding; instead, it likely encouraged more speculative guessing in targeted contexts. This outcome reinforces our central hypothesis stated in Section 7.3 that the model’s extraction is guided by statistical patterns rather than a semantic comprehension of the text’s function.
7.5 Behavior on Non-Latin Pages
@@
Model Details Page Token Model Variant Size Mode Neg. Recall FP Rate Lingua - - T 27.69 2.64 Qwen2.5 VL 32B I+T 65.38 2.66 VL 32B I 73.08 2.09 - 32B T 96.15 0.17 DeepSeek-R1 - 671B T 77.69 0.43
@@
Table 3: Analysis on non-Latin pages: Negative Recall (in pages) and False Positive Rate (averaged percentage of tokens falsely identified as Latin on the whole page).
To gain further insight into model behavior on pages without Latin, we analyzed negative class recall (for page-level detection) and false positive token rates (for segment extraction) as two additional metrics, shown in Table 3. The analysis reveals that most models tend to over-detect Latin. A notable exception is Qwen2.5 with text-only input, which correctly identifies 96% of non-Latin pages, likely due to a stronger sensitivity to linguistic context. In contrast, the statistical baseline Lingua performs poorly, misclassifying over 70% of these pages, which would be a significant issue for large-scale processing. However, a promising finding is that even on misclassified pages, the number of erroneously extracted Latin tokens by the LLMs is generally small, suggesting that simple downstream filtering could effectively mitigate this over-detection problem.
@@ 8 Conclusion
This paper introduced and benchmarked the novel task of zero-shot Latin discovery in historical documents. Our analysis demonstrates that this task is solvable with excellent performance of LLMs without task-specific fine-tuning. However, our key finding is that this success appears to stem not from deep functional semantic understanding, but from the models’ ability to leverage more superficial statistical cues like vocabulary and text patterns, a behavior similar to traditional methods. This may limit the current models’ approach to more nuanced historical tasks.
A particular implication of our work is the high performance of image-only models (up to 99% page-level F1), which enables an efficient two-stage approach for processing vast non-OCR archives. Building on these, our future work will focus on applying our pipeline to the entire ECCO collection and extending the methodology to other corpora, such as the French BnF’s collection.
@@
All of our work is performed on a single collection. Results from this single corpus (18th-century British ECCO texts) should be interpreted with caution, as potential corpus-specific biases may affect broader generalizability, highlighting the importance of future cross-corpus studies. Since Latin has been widely used across Europe, it would be interesting to validate our methods on other collections, e.g., the main texts written in Romance languages, such as French. Nevertheless, we chose to focus exclusively on ECCO for now, since it is the center of our future research interests, which are already straining our computing capacities.
The current work does not utilize a separate validation set, a decision necessitated by the considerable difficulty and cost of annotating the historical data. Consequently, model selection was performed without an independent dataset for optimization before final testing, instead referring to the results from previous work. Incorporating a validation set is an important consideration for future extensions of this research.
All reported results are from single experimental runs in a certain configuration with the maximum level of output determinism and reproducibility. This approach was a necessary compromise due to the significant computational costs and GPU resource constraints associated with evaluating the diverse range of large-scale models investigated. Future work could explore the impact of running variability by conducting multiple trials and tuning seeds, temperatures, and other hyperparameters.
A key direction for future work is to compare the performance of the off-the-shelf models benchmarked in this study against models specifically trained or fine-tuned for the Latin extraction task. While our results demonstrate the surprisingly strong zero-shot capabilities of general-purpose LLMs, fine-tuning would likely establish a much higher state-of-the-art performance ceiling. Moreover, such experiments would be invaluable for identifying the most persistent challenges of the task that remain even after direct, task-specific training, perhaps related to semantic functional understanding of Latin, thereby providing crucial insights for future model development. Our newly created dataset provides the first dedicated resource to enable this line of inquiry.

### v3

@@ Abstract
This paper presents a novel task of extracting low-resourced and noisy Latin fragments from mixed-language historical documents with varied layouts. We benchmark and evaluate the performance of large foundation models against a multimodal dataset of 724 annotated pages. The results demonstrate that reliable Latin detection with contemporary zero-shot models is achievable, yet these models lack a functional comprehension of Latin. This study establishes a comprehensive baseline for processing Latin within mixed-language corpora, supporting quantitative analysis in intellectual history and historical linguistics. Both the dataset and code are available at https://github.com/COMHIS/EACL26-detect-latin.
@@ 2 Problem Definition
Formally, given a document page $D$, let $I_{D}$ denote its image and $T_{D}$ denote its OCR-processed text. A system must perform the following two subtasks:
@@
We structure our problem into two tasks for a more comprehensive evaluation. The core challenge, Task 2, is the fine-grained extraction of Latin text segments, evaluated using per-page token precision and recall. Since these metrics inadequately handle pages with no Latin, we introduce Task 1, a page-level binary classification, to assess performance on these non-Latin instances specifically. We require extracted segments as strings rather than image regions, as this output format better aligns with the capabilities of most MLLMs and enables a simpler, more direct comparison across different input modalities. Detailed definitions of our metrics are provided in Section 5.
@@ 3 Related work
Given its historical importance, Latin has attracted considerable attention within the NLP community (e.g., Sprugnoli et al., 2024; Schulz and Keller, 2016; Gorovaia et al., 2024; Perrone et al., 2021; Burns et al., 2021), though much of this research has centered on small, clean corpora of ancient literary texts. While some recent studies have ventured into Early Modern mixed-language documents (Stüssi and Ströbel, 2024; Volk et al., 2024), these also predominantly rely on manually curated and annotated data. In contrast, our work focuses on the foundational task of Latin discovery: detecting Latin within extensive, unedited, and noisy digitized collections like ECCO Tolonen et al. (2022). This computational approach aims to detect Latin in a vast corpora, while the identified fragments can subsequently be analyzed using a range of established NLP tools developed for classical languages (Johnson et al., 2021; Burns, 2023; Straka and Straková, 2020; Kupari et al., 2024).
@@ 4.1 Sampling and Annotations
These pages were then manually annotated. Annotators were tasked with drawing bounding boxes around all Latin fragments on the page images (see example in Figure 1). These visual annotations are later reliably mapped to text offsets (locations within a string) using ECCO’s OCR positional data for ground truth text extraction. Our annotation guidelines stated marking all instances of Latin text semantically used as Latin. This included single Latin words if presented with explanations in a dictionary, as well as Latin found in headlines, editorial annotations, or footnotes.
The annotation environment was Label Studio Tkachenko et al. (2020-2025). The primary annotation was performed by three scholars familiar with Latin. Following this, an expert in historical texts meticulously reviewed and validated all annotations to ensure accuracy and consistency. During the process, the annotators veto the pages with incompletely OCR-transcribed regions.
@@ 12. Side-note 8
Table 1: Page counts by Latin segment category.
@@ 5.1 OCR Post-Correction and Normalization
Even after the post-correction, residual noise and other variation still remain in the data. Thus, for token-based evaluation, we apply a more traditional rule-based preprocessing pipeline to both predicted and reference strings. This deterministic pipeline, informed by our extensive experience with OCR data and domain-specific knowledge, targets common superficial textual variations and ensures a fair alignment. The pipeline includes Unicode normalization, ligature replacement, lowercasing, digit removal, de-hyphenation, and punctuation stripping. Subsequent to these cleaning operations, the strings are tokenized into word-level units. More details on the processing steps are presented in Appendix B.1.
@@ 5.2 Metrics
The goal of Task 1 is to detect whether a page has Latin on it. We measure this by reporting precision, recall, and F1 score in percentage, along with the F1 score for non-Latin pages to ensure balanced evaluation. To evaluate the Latin segment extraction performance in Task 2, we calculate precision, recall, and F1 score in percentage based on token-level matches between model predictions and the ground truth of the page. A fuzzy matching mechanism is applied to pair predicted and reference tokens one by one. A match is considered valid if the token-level edit distance is not larger than a tunable threshold proportion $\theta$ compared to the ground-truth token length. This approach provides a more flexible and robust evaluation than exact token matching by tolerating minor textual differences at the token level, such as lexical variations and OCR-induced distortions. The pseudocode of the fuzzy matching algorithm is shown in the Appendix B.2. Overall metrics are averaged across the full evaluation set.
@@ 7.1 Experiment Setup
Model Details Page-level (Task 1) Token-level (Task 2) Model Variant Size Mode F1 Precision Recall NL. F1 F1 Precision Recall Lingua - - T 92.58 86.32 99.83 43.11 77.14 77.31 80.07 GPT-4.1 - - I+T 96.03 92.51 99.83 77.00 84.59 88.49 83.54 DeepSeek-R1 - 671B T 97.53 95.34 99.83 87.07 87.00 89.69 86.17 Distill-Llama 70B T 96.18 92.94 99.66 78.34 84.26 87.81 83.62 Distill-Qwen 32B T 95.35 92.55 98.32 74.44 82.84 86.14 82.66 Qwen3 MoE-A22B 235B T 96.90 94.13 99.83 80.52 86.54 89.16 86.00 MoE-A3B 30B T 95.90 93.45 98.48 78.07 84.29 87.29 83.73 - 32B T 95.39 93.25 97.64 75.86 84.07 86.46 83.98 - 14B T 96.61 94.96 98.32 82.85 84.78 88.10 83.71 - 4B T 88.45 87.36 89.56 43.27 75.38 76.30 77.20 Qwen2.5 VL 72B I+T 98.58 98.16 98.99 93.33 83.87 86.98 82.74 VL 72B I 98.82 98.66 98.99 94.57 80.00 82.36 79.72 - 72B T 97.85 100.00 95.79 91.23 80.95 85.31 79.33 VL 32B I+T 96.18 92.94 99.66 78.34 84.32 86.90 83.99 VL 32B I 96.80 94.40 99.33 82.97 79.82 82.64 79.18 - 32B T 97.26 99.13 95.45 88.65 81.57 84.46 80.99 VL 7B I+T 88.36 82.89 94.61 15.91 72.41 78.62 72.23 VL 7B I 95.75 96.74 94.78 81.62 71.42 76.99 70.39 - 7B T 91.86 92.49 91.25 64.18 60.52 65.78 64.57 InternVL3 - 38B I+T 95.62 92.32 99.16 75.00 84.24 84.43 86.67 - 38B I 89.51 87.22 91.92 43.86 56.53 61.11 55.61 - 14B I+T 94.28 90.42 98.48 65.70 81.36 82.63 83.87 - 14B I 90.16 87.86 92.59 47.37 53.51 56.46 55.08 - 8B I+T 90.54 86.39 95.12 41.00 69.04 65.87 82.10 - 8B I 90.82 88.50 93.27 50.88 60.22 60.30 65.63 Gemma3 - 27B I+T 90.77 83.57 99.33 18.92 82.50 84.00 84.79 - 27B I 88.42 82.60 95.12 12.94 60.03 62.78 61.68 - 27B T 94.22 90.03 98.82 64.36 83.79 86.09 84.20
Table 2: Experimental results on selected LLMs, compared with Lingua baseline. “NL. F1” denotes the F1 score for identifying Non-Latin pages. “I” and “T” indicate image and text input separately.
7.2 Model Results and Analysis
@@
Finally, we address the issue of robustness in rejecting non-Latin pages. Beyond standard detection metrics, the NL. F1 column reveals a divergence in model behavior: while detection recall is generally high, the ability to correctly reject non-Latin pages varies. This instability is most pronounced in Gemma3 with image-only input. It corroborates our earlier observation regarding its weak visual grounding. Yet, this issue extends beyond a single model, and we analyze further in Section 7.3.
7.3 Behavior on Non-Latin Pages
To further quantify the over-sensitivity observed in the main results, we analyzed non-Latin class recall (for page-level detection) and false positive token rates (for false Latin segment extraction) as two additional metrics, shown in Table 3. The analysis reveals that most models tend to over-detect Latin, particularly the smaller ones. The statistical baseline Lingua performs poorly, misclassifying over 70% of these pages, which would be a significant issue for large-scale processing.
More critically, multimodal inputs can overwhelm smaller models: Qwen2.5-VL-7B suffers a catastrophic collapse, indicating that overlapped visual and textual signals cause the model to hallucinate Latin across half the page. Crucially, this is a failure of fusion rather than perception: when fed with only images, the same model recovers much of its rejection capability, confirming that long and noisy OCR text acts as a distractor that overrides visual grounding in smaller architectures.
In contrast, the most robust performance comes from Qwen2.5-32B with text-only input, likely due to a stronger sensitivity to linguistic context provided by model scale. However, a promising finding is that even on misclassified pages, the number of erroneously extracted Latin tokens by the LLMs is generally small, suggesting that simple downstream filtering could effectively mitigate this over-detection problem.
@@ 7.4 Performance by Category
Model Details Page Token Model Variant Size Mode NL. Recall $\uparrow$ FP Rate $\downarrow$ Lingua - - T 27.69 2.64 DeepSeek-R1 - 671B T 77.69 0.43 Qwen3 - 4B T 40.77 4.97 Qwen2.5 VL 32B I+T 65.38 2.66 VL 32B I 73.08 2.09 - 32B T 96.15 0.17 VL 7B I+T 10.77 55.54 VL 7B I 85.38 2.97 - 7B T 66.15 10.07
@@
Table 3: Analysis on non-Latin pages: Non-Latin Recall in pages and False Positive Rate showing averaged percentage of tokens falsely identified as Latin on truly non‑Latin pages.
7.5 Impact of Prompt Variations
To assess the robustness of our findings, we first evaluated the performance stability of four representative models across six distinct prompt strategies, as summarized in Table 4. We observe a consistent insensitivity to instruction phrasing across diverse architectures, with minimal Token F1 variance ($\sigma<1.1$) and no consistent gains, for both text-only and multimodal models. This statistical evidence suggests that extraction capabilities are largely anchored by the models’ intrinsic domain knowledge rather than specific prompt engineering.
Focusing on the detailed behavior of the top-performing Qwen2.5-VL-32B model with multimodal input (shown in Figure 3), most strategies result in only marginal changes, typically a trade-off between precision and recall, or different tasks, rather than a definitive improvement. We highlight the most salient strategies here, with full tables of all the prompts and results available in Appendix E.2. For instance, instructing the model to include “single-word” segments lowered page-level precision by encouraging over-extraction on short segments, while negative constraints like “no borrow” slightly increased token-level precision through a more conservative behavior by excluding the borrowed Latin words in other languages. Most revealingly, the “partial categories” strategy, directing the model to focus on its known weak areas, increased token recall but at the cost of a drop in precision. This indicates the prompt did not grant the model a deeper functional understanding; instead, it likely encouraged more speculative guessing in targeted contexts. This outcome reinforces our central hypothesis stated in Section 7.4 that the model’s extraction is guided by statistical patterns rather than a semantic comprehension of the text’s function.
Model Variant Size Mode Page F1 Tok. F1 DeepSeek-R1 Distill 70B T $96.28_{\pm 1.06}$ $84.27_{\pm 0.20}$ Qwen3 - 14B T $96.47_{\pm 1.00}$ $84.42_{\pm 0.62}$ Qwen2.5 - 32B T $97.14_{\pm 0.75}$ $81.89_{\pm 1.06}$ Qwen2.5 VL 32B I+T $95.58_{\pm 1.46}$ $84.44_{\pm 0.55}$ VL 32B I $95.72_{\pm 1.79}$ $79.51_{\pm 0.79}$
Table 4: Prompt robustness statistics: performance reported as $Mean_{\pm Std}$ across 6 prompt strategies.
Figure 3: Impact of prompting on Qwen2.5-VL-32B.
7.6 Impact of OCR Quality
As detailed in Section 5.1, evaluating models on the ECCO corpus is complicated by significant discrepancies between legacy OCR and modern visual processing. To ensure meaningful assessment across all model types, we employ an OpenAI o1-based post-correction pipeline Kanerva et al. (2025).
To quantify the necessity of this normalization, we conducted an ablation study using raw, uncorrected OCR transcripts, shown as Table 5. We observe that raw OCR creates a noise barrier that affects all models utilizing text input (including Vision-Language models fed with OCR text), with pure text models suffering the most severe degradation (>14>14 point drop). Consequently, the performance degrades so severely that text-dependent models fall behind the best image-only model. This confirms that raw ECCO OCR constructs an excessive noise barrier that surpasses the detection task itself. Thus, our post-correction pipeline is not an o1-assisted “shortcut” for OCR-dependent models, but a necessary normalization to ensure fair cross-modal comparison.
Model & Input Page F1 Tok. F1 $\Delta$ Tok. F1 Text-only Models DeepSeek-R1 96.65 72.26 -14.74 w/ Cleaned OCR 97.53 87.00 Qwen3-14B 93.33 67.12 -17.66 w/ Cleaned OCR 96.61 84.78 Qwen2.5-32B 96.02 67.48 -14.09 w/ Cleaned OCR 97.26 81.57 Vision-Language Models Qwen2.5-VL-32B 95.38 79.24 -5.08 w/ Cleaned OCR 96.18 84.32 InternVL3-38B 93.74 74.56 -9.68 w/ Cleaned OCR 95.62 84.24 Image-only Models Qwen2.5-VL-32B (I) 96.80 79.82 - InternVL3-38B (I) 89.51 56.53 -
Table 5: Ablation on OCR quality. $\Delta$ Tok. F1 denotes the token-level F1 change compared to the models with cleaned OCR text input.
@@ 8 Conclusion
This paper introduced and benchmarked the novel task of zero-shot Latin discovery in historical documents. Our analysis demonstrates that this task is solvable with excellent performance of LLMs without task-specific fine-tuning. However, our key finding is that this success appears to stem not from deep functional semantic understanding, but from the models’ ability to leverage more superficial statistical cues like vocabulary and text patterns, a behavior similar to traditional methods. This may limit the current models’ approach to more nuanced historical tasks, suggesting that foundation models cannot replace human interpretative expertise.
A particular implication of our work is the high performance of image-only models (up to approximately 99% page-level F1), which enables an efficient two-stage approach for processing vast non-OCR archives. This capability may effectively unlock the “dark archives” of digitized heritage. Building on these, our future work will focus on applying our pipeline to the entire ECCO collection and extending the methodology to other corpora, such as the French BnF’s collection.
@@
Our findings rely on a single corpus (18th-century British ECCO). While this aligns with our research scope, corpus-specific biases may limit generalizability, necessitating future validation on diverse collections, such as Romance-language texts. Additionally, due to the high cost of annotation, we lacked a separate validation set for hyperparameter tuning. Model selection relied on literature baselines rather than independent optimization.
Computational constraints limited us to single experimental runs with deterministic settings. While ensuring reproducibility, this approach does not capture the potential variance inherent in stochastic generation. Finally, this study benchmarks off-the-shelf models in a zero-shot setting. We acknowledge that task-specific fine-tuning would likely yield higher performance ceilings. Moreover, such experiments would be invaluable for identifying the most persistent challenges of the task that remain even after direct, task-specific training—perhaps related to deep functional semantic understanding of Latin—thereby providing crucial insights for future model development. Our newly created dataset provides the first dedicated resource to enable this line of inquiry.